OpenStudy (anonymous):

Use the functions f(x) = 5x − 12 and g(x) = 11x + 2 to complete the function operations listed below. Part A: Find (f + g)(x). Show your work. (3 points) Part B: Find (f ⋅ g)(x). Show your work. (3 points) Part C: Find f[g(x)]. Show your work. (4 points)

OpenStudy (anonymous):

@saseal

OpenStudy (anonymous):

\[(f+g)(x)=5x-12+11x+2\]\[(f \times g)(x) = (5x-12)(11x+2)\]\[f[g(x)]=5(11x+2)-12\]

OpenStudy (anonymous):

is this part a,b, and c

OpenStudy (anonymous):

@saseal

OpenStudy (anonymous):

@saseal

OpenStudy (anonymous):

@saseal

OpenStudy (anonymous):

they are for all the parts

OpenStudy (anonymous):

so this is the answer with the work?

OpenStudy (anonymous):

no you need to simplify it

OpenStudy (anonymous):

can u just tell me the answer with the work plz

OpenStudy (anonymous):

@AnimalLover8

OpenStudy (anonymous):

?

OpenStudy (animallover8):

One minute. I'm kinda' busy.